# Probing Quantum Anomalous Hall States in Twisted Bilayer WSe2 via Attractive Polaron Spectroscopy

### v1

@@ Conclusion
Looking ahead, several exciting research avenues can be built upon our findings. One promising direction is the exploration of fractional Chern insulators [33], which have recently been observed in tMoTe2 [1, 2, 3, 4, 34, 35, 36], and may be realized in tWSe2 given its favorable material properties. Moreover, the presence of different Chern numbers in higher moire bands compared to MoTe2 may lead to the realization of other correlated states  [37].
More intriguing is the realization of recently predicted topological excitons [38], accessible in a modified device architecture designed to host long-lived interlayer excitons. The coexistence of strong exciton-exciton interactions [39, 17, 16, 40] and tunable excitonic band topology [38] offers a compelling pathway toward the realization of bosonic fractional Chern insulators [41, 42, 43]. Such long-sought phase are bosonic analogue of fermionic fractional Chern insulators [44, 45, 46, 47, 48], which has remained unobserved in any system to date. More generally, this system could be considered as a promising candidate for optical probe and manipulation of strongly-correlated electron-photon systems [49, 50]. Specifically, interesting prospects are optical detection and manipulation of topological states [51, 52].
.0.1 Methods
Device fabrication:
@@
Optical measurements:
Optical measurements were performed using confocal microscopy in an optically accessible dilution fridge (BlueFors) with vertical magnetic fields up to 9T and temperatures down to 60 mK [53]. A halogen lamp was used as the light source. The input light was spectrally filtered to 650 to 900 nm range. A low-temperature microscope objective was used to focus the light onto the sample. The light intensity on the sample was kept around 0.14 nW/$\mu$m2 to minimize its effects on the electronic states. The reflected light was collected by the same objective and analyzed by a spectrometer equipped with a charge coupled device array to obtain spectrum. The RC spectrum is defined as $(R-R_{0})/R$, where $R_{0}$ the reference spectrum was taken for the sample at a high doping density with quenched excitonic resonances. For all the zero magnetic field data presented in Fig. 2 and  4, we polarize the system with a positive vertical magnetic field and then ramp back to perform the zero magnetic field measurement. The RMCD was used to study the magnetic properties of the samples. A combination of a linear polarizer and a quarter-wave plate was used to generate circularly polarized light ($\sigma_{-}$ and $\sigma_{+}$) on the sample. The RMCD spectrum is defined as $\frac{\sigma_{-}-\sigma_{+}}{{\sigma_{-}+\sigma_{+}}}$.
@@
Twisted Angle Calibration:
We calibrate the twisted angle of the tWSe2 presented in the main text using the quantum oscillations observed optically under an out-of-plane magnetic field of 8.8T. In Extended Data Fig. 2, we present the doping dependent $RC$ response of a monolayer tWSe2 under $\sigma_{+}$ excitation. The oscillation signal formed as a result of the formation of the Landau levels (LLs). The LL period is determined to be 0.244 V for the single gate response, from which we deduce a hole density change of 1.75×$\times 10^{12}$ cm-2 per volt for data presented in the main text. Combining with the moiré period $\Delta V_{M}$ extracted from the doping dependent $RC$ data, we obtain a moiré density $n_{M}$ of 1.31×$\times 10^{12}$ cm-2. The twisted angle calibration is further corroborated by the doping density calibration, as described in the next section.
Doping density calibration:
@@
.0.2 Data availability
All of the data that support the findings of this study are reported in the main text and Supplementary Information. Source data are available from the corresponding authors on reasonable request.
.0.3 Acknowledgements
We thank Ataç İmamoğlu, Liang Fu, Sankar Das Sarma, Ajit Srivastava, Patrick Knüppel, Zhengguang Lu, Xiaowei Zhang for fruitful discussions.
.0.4 Competing interests
The authors declare no competing interests.
.0.5 Contributions
B.G., Y.Z., and M.H. conceived the project. B.G. performed the optical measurements and analyzed the data. M.X. did the theoretical calculation. M.X. and T.S.H. contributed to the theoretical understanding. L.Z. and B.G. fabricated the device. L.Z., M.G., D.S., P.U., G.A., R.M., M.J.M, and S.S. assisted with the experiment. S.J. and H.J. exfoliated the TMD flakes. T.T. and K.W. provided the bulk h-BN crystals. All authors contributed the writing of the manuscript. Y.Z. and M.H. supervised the work. All authors discussed the results and contributed to the manuscript.

### v2

@@ Conclusion
Looking ahead, several exciting research avenues can be built upon our findings. One promising direction is the exploration of fractional Chern insulators [9], which have recently been observed in tMoTe2 [6, 54, 29, 50, 7, 42, 3], and whose realization in tWSe2 would open access to new correlated topological phases. Furthermore, similar to the recently demonstrated coexistence of QAH and superconductivity in tMoTe2 [51], we anticipate intriguing interplay between topological states and superconductivity in tWSe2[13, 45]. Moreover, the presence of different Chern numbers in higher moire bands compared to MoTe2 may lead to the realization of other correlated states  [55]. More intriguing is the realization of recently predicted topological excitons [46], accessible in a modified device architecture designed to host long-lived interlayer excitons. The coexistence of strong exciton-exciton interactions [30, 12, 48, 20] and tunable excitonic band topology [46] offers a compelling pathway toward the realization of bosonic fractional Chern insulators [14, 22, 41]. Such long-sought phase are bosonic analogue of fermionic fractional Chern insulators [33, 25, 34, 37, 36], which has remained unobserved in any system to date. More generally, this system could be considered as a promising candidate for optical probe and manipulation of strongly-correlated electron-photon systems [5, 4]. Specifically, interesting prospects are optical detection and manipulation of topological states [21, 24].
Methods
Device fabrication
@@ Methods
Optical measurements
[…]
@@
Twist Angle Calibration
[…]
Doping density calibration
@@
Extraction and analysis of the Streda signal
To extract the Streda response, we first obtain polarization-resolved reflection-contrast (RC) spectra as a function of gate voltage and magnetic field. For each magnetic-field value, an energy window containing the attractive-polaron (AP) resonance is integrated to yield an intensity map $I_{\rm{AP}}(\nu,\mu_{0}H)$, where $\nu$ is the hole filling. The resulting $I_{\rm{AP}}(\nu)$ curve exhibits local maxima at incompressible fillings. We identify this maximum from the derivative of $I_{\rm{AP}}$ with respect to $\nu$ and treat its position $\nu_{\rm{max}}(B)$ as the quantity entering the Streda analysis.
The uncertainty in each extracted $\nu_{\rm{max}}$ is estimated by computing the centroid of the AP peak at fixed magnetic field and defining an error range corresponding to a deviation of the integrated intensity from the centroid. For the sharp $\nu=1$ feature we use a 3% deviation; for the broader $\nu=3$ feature we use 5%.
The Streda slope $d\nu/dB$ is then obtained from a linear fit to $\nu_{\rm{max}}(B)$ near the filling of interest. Using the Streda relation $C=\frac{h}{e}\frac{d\nu}{dB}$, we extract the corresponding Chern number. To verify robustness, we repeat the analysis using several different integration windows for the AP resonance (widths of 3.05 meV, 5.71 meV, and 8.85 meV). All choices yield consistent slopes within the extracted uncertainties.
Acknowledgements: We thank Ataç İmamoğlu, Liang Fu, Sankar Das Sarma, Ajit Srivastava, Patrick Knüppel, Xiaowei Zhang for fruitful discussions.
Author Contributions: B.G., Y.Z., and M.H. conceived the project. B.G. performed the optical measurements and analyzed the data. M.X. did the theoretical calculation. M.X. and T.S.H. contributed to the theoretical understanding. L.Z. and B.G. fabricated the device. L.Z., M.G., D.S., P.U., G.A., R.M., M.J.M, and S.S. assisted with the experiment. S.J. and H.J. exfoliated the TMD flakes. T.T. and K.W. provided the bulk h-BN crystals. All authors contributed the writing of the manuscript. Y.Z. and M.H. supervised the work. All authors discussed the results and contributed to the manuscript.
Competing interests: The authors declare no competing interests.
Data availability: All of the data that support the findings of this study are reported in the main text and Supplementary Information. Source data are available from the corresponding authors on reasonable request.